# Robust Model-Based Reinforcement Learning with an Adversarial Auxiliary Model

### v1

@@ 2.1 Model-Based Reinforcement Learning
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi,P,\rho_{0}}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t},s_{t+1})\mid s_{0}\sim\rho_{0}\right]$
@@ 2.2 Robust Markov Decision Processes
In a traditional MDP, the agent optimizes its policy in a static transition model $P$. However, in some real-world problems, the transition model can change over time. Hence, we can define a Robust MDP (Wiesemann et al., 2013) where the agent acts in an unknown MDP $P\in\mathcal{P}$ that is a sample from an uncertainty set $\mathcal{P}$. The robust objective $J_{\mathcal{P},\pi}$ can now be defined to maximize an objective function in the worst-case MDP of a given uncertainty set. This objective is formally stated in Eq. (1).
@@ 3 Auxiliary Adversarial Model
The goal of this section is to tackle the inner-loop problem of the robust objective, as defined by the minimization problem in Eqn. (1), i.e. approximating the worst-case MDP, denoted as $P^{*}\in\mathcal{P}$, where we choose $\mathcal{P}$ to be the KL uncertainty set centered around the nominal model $\bar{P}$. This choice of uncertainty set follows a common choice in literature (Wang et al., 2024; Hu & Hong, 2013). To describe our methodology, this section first (Sec. 3.1) introduces the auxiliary adversarial model as an addition to traditional world model learning (e.g. via maximum likelihood estimation (Janner et al., 2019)). The auxiliary model has a well-defined KL divergence with the approximated nominal model. Secondly (Sec. 3.2), we introduce the loss function to train the auxiliary model to maintain a low KL divergence with the nominal transition model, whilst also learning to be pessimistic (i.e., minimizing the return of the transition). Finally, we propose Robust MBPO (RMBPO), an algorithm that incorporates the auxiliary model to improve the robustness of the learned policy.
@@ 3.1 Auxiliary Model
Since we work within the context of Model-Based RL, we have direct access to a parameterized approximation, $p_{\theta}(s^{\prime},r|s,a)$, of the nominal transition model $\bar{P}(.)$. However, this does not directly provide us with a method to approximate $D_{KL}(p_{\theta}||\bar{P})$, since we do not have access to the simulated transition probabilities, $\bar{P}(s^{\prime},r|s,a)$, needed to construct the KL uncertainty set. Hence, we propose to not directly try to approximate the pessimistic transition model, thus leaving $p_{\theta}$ untouched. As an alternative, we propose an auxiliary parameterized model, $g_{\psi}$, which takes as input the outputs of the learned transition model $p_{\theta}$, in addition to $s$ and $a$. Next states and rewards can now be sampled according to Eqn. (3).
@@ 3.2 Training the Auxiliary Model
The goal of the auxiliary model is to minimize the value under the current policy while remaining within the desired uncertainty set $\mathcal{P}_{KL}$. To incentivize both these objectives, the auxiliary model is trained by minimizing a sum of two loss functions, displayed in Eqn. (4). The first term is the Kullback-Leibler (KL) divergence between the normal model $p_{\theta}$ and the auxiliary model $g_{\psi}$, which ensures that the auxiliary model does not deviate too far from the (approximated) dynamics of the nominal environment. Therefore, this part of the loss function incentivizes the model to remain within the KL uncertainty set $\mathcal{P}_{KL}$. The second term of the loss function minimizes the return of the current state transition, as proposed by Rigter et al. (2022). By lowering the (log) likelihood of transitions with a high return and heightening the likelihood of transitions with a low return, the auxiliary model will become a pessimistic approximation of $p_{\theta}$ and $\bar{P}$.
@@
In Eqn. (4), sg() defines the stop_grad operator, no gradient is computed for the value, only for $g_{\psi}$. $V(.)$ denotes the (approximate) value function, used to estimate the expected return after the transition. The hyperparameter $\eta$ controls the influence of the value function: for a small $\eta$, $p_{\theta}\approx g_{\psi}$ and therefore $\mathcal{P}_{KL}$ is small, for larger values of $\eta$, $g_{\psi}$ will grow more pessimistic and therefore $\mathcal{P}_{KL}$ can be large. The values of $\eta$ that were used in this work can be found in Appendix A.
@@ 3.3 Proposed Algorithm
To improve policy robustness, we combine the auxiliary model with MBPO (Janner et al., 2019) to become RMBPO. MBPO approximates the training environment by maximizing the likelihood of experienced transitions under its learned model $p_{\theta}$. This model is a neural network that predicts a mean and covariance matrix over the next states and rewards, conditioned on the current state and action. On-policy rollouts are then performed on the learned model. Finally, the unrolled data is used to update a policy via Soft-Actor Critic (SAC) (Haarnoja et al., 2018). We modify MBPO by training an auxiliary model in addition to the existing model, via Eqn (4). Since these two models are trained separately, the auxiliary model learning does not hinder the accuracy or precision of $p_{\theta}$. During the model unroll, we pass the current state through the learned model $p_{\theta}$, after which we use the output of that model ($\mu_{\theta},\Sigma_{\theta})$ as input to the auxiliary model. The auxiliary model will then predict a modified ($\mu_{\psi},\Sigma_{\psi})$ as an approximation to the worst-case transition model in $\mathcal{P}_{KL}$. Relating to Sec. 2.2, the auxiliary model tries to solve the inner-loop problem, while SAC tries to maximize the outer-loop problem. These two components act as two players in a zero-sum Markov game. This algorithm is fully described in Algorithm 1, where our additions are highlighted in blue. Following other works (Wang et al., 2024; Zhou et al., 2024), we add a small amount of action noise to the environment, otherwise, the uncertainty set would not be well-defined. More details on the action noise are provided in Appendix B.
@@
8:         Update model parameters $\psi$ according to Eqn. (4): $\psi\leftarrow\psi-\lambda_{a}\hat{\nabla}_{\psi}J_{g}(\psi,\mathcal{D}_{env},p_{\theta},\pi_{\phi}$)
@@
14:              On-policy rollout according to Eqn. (3) starting from $s_{t}$ using policy $\pi_{\phi}$; add to $\mathcal{D}_{model}$
@@ 4 Results
The following section aims to answer two main research questions: "Does the auxiliary model learn pessimistic state transitions?" and "Can the auxiliary model make a learned policy more robust?". The first question is investigated in Sec 4.2, where we perform a case study on the Hopper-v3 environment to examine which changes are made by the auxiliary model. The second question is investigated in 4.1, where we plot the performance of RMBPO compared to MBPO under distorted evaluation environments.
@@ Appendix B Implementation details
Following related work (Zhou et al., 2024), we slightly modify the mentioned MuJoCo environments by adding Gaussian noise to the action: $a_{t}\leftarrow a_{t}+\mathcal{N}(0,5e-3)$. Only in Pendulum-v2, we use a significantly higher noise variance of 0.2. Since this action noise is invisible to the agent, it introduces stochasticity in the MDP. Inspired by the existing MBPO world model, we standardize the outputs of $p_{\theta}$ before providing them as inputs to $g_{\psi}$, this showed incremental stability improvements in some environments. As proposed in appendix A.1 by Rigter et al. (2022), we subtract $V^{\theta,\psi}_{\phi}(s)$ as a baseline from the return in Eqn. 4, this does not influence the expectation of the gradient but significantly reduces its variance. Note that MBPO does not employ a value network directly, but this does not pose an issue since on-policy samples from the Q-network will provide the same expected gradient, given a large enough minibatch size.

### v2

@@ 2.1 Model-Based Reinforcement Learning
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi,P,\rho_{0}}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t},s_{t+1})\mid s_{0}\sim\rho_{0}\right]$ (1)
@@ 2.2 Robust Markov Decision Processes
In a traditional MDP, the agent optimizes its policy in a static transition model $P$. However, in some real-world problems, the transition model can change over time. Hence, we can define a Robust MDP (Wiesemann et al., 2013) where the agent acts in an unknown MDP $P\in\mathcal{P}$ that is a sample from an uncertainty set $\mathcal{P}$. The robust objective $J_{\mathcal{P},\pi}$ can now be defined to maximize an objective function in the worst-case MDP of a given uncertainty set. This objective is formally stated in Eq. 2.
@@ 3 Auxiliary Adversarial Model
The goal of this section is to tackle the inner-loop problem of the robust objective, as defined by the minimization problem in Eq. 2, i.e. approximating the worst-case MDP, denoted as $P^{*}\in\mathcal{P}$, where we choose $\mathcal{P}$ to be the KL uncertainty set centered around the nominal model $\bar{P}$. This choice of uncertainty set follows a common choice in literature (Wang et al., 2024; Hu & Hong, 2013). To describe our methodology, this section first (Section 3.1) introduces the auxiliary adversarial model as an addition to traditional world model learning (e.g. via maximum likelihood estimation (Janner et al., 2019)). The auxiliary model has a well-defined KL divergence with the approximated nominal model. Secondly (Section 3.2), we introduce the loss function to train the auxiliary model to maintain a low KL divergence with the nominal transition model, whilst also learning to be pessimistic (i.e., minimizing the return of the transition). Finally, we propose Robust MBPO (RMBPO), an algorithm that incorporates the auxiliary model to improve the robustness of the learned policy.
@@ 3.1 Auxiliary Model
Since we work within the context of Model-Based RL, we have direct access to a parameterized approximation, $p_{\theta}(s^{\prime},r|s,a)$, of the nominal transition model $\bar{P}(.)$. However, this does not directly provide us with a method to approximate $D_{KL}(p_{\theta}||\bar{P})$, since we do not have access to the simulated transition probabilities, $\bar{P}(s^{\prime},r|s,a)$, needed to construct the KL uncertainty set. Hence, we propose to not directly try to approximate the pessimistic transition model, thus leaving $p_{\theta}$ untouched. As an alternative, we propose an auxiliary parameterized model, $g_{\psi}$, which takes as input the outputs of the learned transition model $p_{\theta}$, in addition to $s$ and $a$. Next states and rewards can now be sampled according to Eq. 4.
@@ 3.2 Training the Auxiliary Model
The goal of the auxiliary model is to minimize the value under the current policy while remaining within the desired uncertainty set $\mathcal{P}_{KL}$. To incentivize both these objectives, the auxiliary model is trained by minimizing a sum of two loss functions, displayed in Eq. 5. The first term is the Kullback-Leibler (KL) divergence between the normal model $p_{\theta}$ and the auxiliary model $g_{\psi}$, which ensures that the auxiliary model does not deviate too far from the (approximated) dynamics of the nominal environment. Therefore, this part of the loss function incentivizes the model to remain within the KL uncertainty set $\mathcal{P}_{KL}$. The second term of the loss function minimizes the return of the current state transition, as proposed by Rigter et al. (2022). By lowering the (log) likelihood of transitions with a high return and heightening the likelihood of transitions with a low return, the auxiliary model will become a pessimistic approximation of $p_{\theta}$ and $\bar{P}$.
@@
In Eq. 5, sg() defines the stop_grad operator, no gradient is computed for the value, only for $g_{\psi}$. $V(.)$ denotes the (approximate) value function, used to estimate the expected return after the transition. The hyperparameter $\eta$ controls the influence of the value function: for a small $\eta$, $p_{\theta}\approx g_{\psi}$ and therefore $\mathcal{P}_{KL}$ is small, for larger values of $\eta$, $g_{\psi}$ will grow more pessimistic and therefore $\mathcal{P}_{KL}$ can be large. The values of $\eta$ that were used in this work can be found in Appendix A.
@@ 3.3 Proposed Algorithm
To improve policy robustness, we combine the auxiliary model with MBPO (Janner et al., 2019) to become RMBPO. MBPO approximates the training environment by maximizing the likelihood of experienced transitions under its learned model $p_{\theta}$. This model is a neural network that predicts a mean and covariance matrix over the next states and rewards, conditioned on the current state and action. On-policy rollouts are then performed on the learned model. Finally, the unrolled data is used to update a policy via Soft-Actor Critic (SAC) (Haarnoja et al., 2018). We modify MBPO by training an auxiliary model in addition to the existing model, via Eq. 5. Since these two models are trained separately, the auxiliary model learning does not hinder the accuracy or precision of $p_{\theta}$. During the model unroll, we pass the current state through the learned model $p_{\theta}$, after which we use the output of that model ($\mu_{\theta},\Sigma_{\theta})$ as input to the auxiliary model. The auxiliary model will then predict a modified ($\mu_{\psi},\Sigma_{\psi})$ as an approximation to the worst-case transition model in $\mathcal{P}_{KL}$. Relating to Section 2.2, the auxiliary model tries to solve the inner-loop problem, while SAC tries to maximize the outer-loop problem. These two components act as two players in a zero-sum Markov game. This algorithm is fully described in Algorithm 1, where our additions are highlighted in blue. Following other works (Wang et al., 2024; Zhou et al., 2024), we add a small amount of action noise to the environment, otherwise, the uncertainty set would not be well-defined. More details on the action noise are provided in Appendix B.
@@
8:         Update model parameters $\psi$ according to Eq. 5: $\psi\leftarrow\psi-\lambda_{a}\hat{\nabla}_{\psi}J_{g}(\psi,\mathcal{D}_{env},p_{\theta},\pi_{\phi}$)
@@
14:              On-policy rollout according to Eq. 4 starting from $s_{t}$ using policy $\pi_{\phi}$; add to $\mathcal{D}_{model}$
@@ 4 Results
The following section aims to answer two main research questions: "Does the auxiliary model learn pessimistic state transitions?" and "Can the auxiliary model make a learned policy more robust?". The first question is investigated in Section 4.2, where we perform a case study on the Hopper-v3 environment to examine which changes are made by the auxiliary model. The second question is investigated in Section 4.1, where we plot the performance of RMBPO compared to MBPO under distorted evaluation environments.
@@ Appendix B Implementation details
Following related work (Zhou et al., 2024), we slightly modify the mentioned MuJoCo environments by adding Gaussian noise to the action: $a_{t}\leftarrow a_{t}+\mathcal{N}(0,5e-3)$. Only in Pendulum-v2, we use a significantly higher noise variance of 0.2. Since this action noise is invisible to the agent, it introduces stochasticity in the MDP. Inspired by the existing MBPO world model, we standardize the outputs of $p_{\theta}$ before providing them as inputs to $g_{\psi}$, this showed incremental stability improvements in some environments. As proposed in appendix A.1 by Rigter et al. (2022), we subtract $V^{\theta,\psi}_{\phi}(s)$ as a baseline from the return in Eq. 5, this does not influence the expectation of the gradient but significantly reduces its variance. Note that MBPO does not employ a value network directly, but this does not pose an issue since on-policy samples from the Q-network will provide the same expected gradient, given a large enough minibatch size.